OpenStudy (anonymous):

Solve the system 2x + 3y = 3 and 3x – 2y = 11 by using graph paper or graphing technology. What is the solution to the system? a. (–3, 3) b.(–1, –7) c. (1, –4) d. (3, –1)

OpenStudy (anonymous):

You should rearrange the formulas into y=mx + b form and the graph them on the same graph. The point where the intersect is the solution to the system. Do you know how to do this?

OpenStudy (anonymous):

If you have a graphing calculator rearrange the formula for y=mx + b and plug into the calculator same deal look for the intersection. You could also solve like this 3/2(2x+3y = 3) is 3x + (9/2)y = 9/2 then subtract second equation from first to get 3x + (9/2)y = 9/2 -3x - 2y = 11 becomes 13/2y = -13/2 y = -1 3(x) - 2(-1) = 11 so x = 3 So you answer should come out to d. (3,-1)

OpenStudy (anonymous):

You could also find an online graphing application that allows you to graph to things simultaneously.

OpenStudy (anonymous):

two things**

OpenStudy (anonymous):

@mathdood (<-nice), can you explain to me a lil more how you got that answer.. like where the fudge monkeys that 9 came from lol

OpenStudy (anonymous):

#so lost